OpenStudy (anonymous):

HELP! what is the magnitude of the vector with an initial point (2,1) and terminal point (-13,9)

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points, or vector magnitude}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ 1}})\quad &({\color{red}{ -13}}\quad ,&{\color{blue}{ 9}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

is the answer 7 ?

OpenStudy (amistre64):

is that a guess? becuase its not close at all

OpenStudy (anonymous):

no i worked out the problem he gave me

OpenStudy (anonymous):

@jdoe0001 can you help me some more

OpenStudy (jdoe0001):

well... anything confusing there? looks pretty straight forward

OpenStudy (anonymous):

is it 7 or am i wrong

OpenStudy (jdoe0001):

well what did you get for \(\bf (x_2-x_1)^2\) ?

OpenStudy (amistre64):

one problem i have with the formula is i simply put the numbers in the wrong places .... i would rather step it out to avoid the error

OpenStudy (anonymous):

(-15)^2

OpenStudy (jdoe0001):

yeap that'd 225, then get the other and sum them up :)

OpenStudy (jdoe0001):

\(\bf \bf \textit{distance between 2 points, or vector magnitude}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ 1}})\quad &({\color{red}{ -13}}\quad ,&{\color{blue}{ 9}}) \end{array}\qquad d = \sqrt{({\color{red}{ -13}}-{\color{red}{ 2}})^2 + ({\color{blue}{ 9}}-{\color{blue}{ 1}})^2} \\ \quad \\ d=\sqrt{(-15)^2+(8)^2}\implies d=\sqrt{225+64}\)

OpenStudy (anonymous):

so its 17

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

can you help me with another problem

OpenStudy (jdoe0001):

easier if you post anew, if we dunno then someone else may know, we can also revise each other, also I'd be dashing in a few mins

OpenStudy (anonymous):

ok